OpenStudy (anonymous):

After a discount of 10%, the price of a pair of jeans, including tax, was $69.62. The following equation can be used to find x, the price of the pair of jeans without the discount. x - 0.1x + 0.07(x - 0.1x) = 69.62 What is the price of the pair of jeans without the discount?

OpenStudy (tanner23456):

first multiply the entire equation (both sides) by 100 to get everything out of decimal form. 100(x-0.1x+0.07(x-0.1x)=69.62) 100x-10x+7(x-0.1x)=6962 90x+7x-0.7x = 6962 you can multiply by ten again to get that last one out of decimal form 10(90x+7x-0.7x = 6962) 900x + 70x - 7x = 69620 963x = 69620 divide the 963 by both sides x = 72.29

OpenStudy (anonymous):

thank you so mach!! that helped a lot!

OpenStudy (tanner23456):

No problem. Glad to help